# Meta-Property Graphs: Extending Property Graphs with Metadata Awareness and Reification

### v1

@@ 5.2 Semantics
Conditions: $\llbracket\mathcal{E}_{1}=\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ =$=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (46) $\llbracket\mathcal{E}_{1}<\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ =$=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}<\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (47) $\llbracket x:\ell\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&\text{if }\llbracket x\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\in\mu(\lambda(\llbracket x\rrbracket^{\beta}_{G}))\\ False&\text{if }\llbracket x\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\notin\mu(\lambda(\llbracket x\rrbracket^{\beta}_{G}))\end{cases}$ (48)
Logical operations: $\llbracket\Phi_{1}\ \text{{AND}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\wedge\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (49) $\llbracket\Phi_{1}\ \text{{OR}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\vee\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (50) $\llbracket\text{{NOT}}\ \Phi\rrbracket^{\beta}_{G}$ =$=\neg\llbracket\Phi\rrbracket^{\beta}_{G}$ (51)
Set operations, where $c\in\mathcal{K}\cup\mathcal{V}\cup\mathcal{L}$: $\llbracket c\ \text{{ELEMENTOF}}\ y\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&c\in\mu(\llbracket y\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (52) $\llbracket\text{{SUBSETEQ}}(x,y)\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&\mu(\llbracket x\rrbracket^{\beta}_{G})\subseteq\mu(\llbracket y\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (53)
@@
Clauses and queries: $\llbracket\text{{MATCH}}\ \Pi\rrbracket_{G}(T)$ $=\bigcup_{\beta\in T}\{\beta\Join\beta^{\prime}\mid\beta^{\prime}\in\llbracket\Pi\rrbracket_{G},\beta\sim\beta^{\prime}\}$ (54) $\llbracket\text{{FILTER}}\ \Phi\rrbracket_{G}(T)$ =$=\{\beta\in T\mid\llbracket\Phi\rrbracket^{\beta}_{G}=True\}$ (55) $\llbracket\mathcal{C}\ \mathcal{Q}\rrbracket_{G}(T)$ $=\llbracket\mathcal{Q}\rrbracket_{G}(\llbracket\mathcal{C}\rrbracket_{G}(T))$ (56) $,\displaystyle\llbracket\text{{RETURN}}\ \mathcal{E}_{1}\ \text{{AS}}\ x_{1},$$\ldots,\mathcal{E}_{n}\ \text{{AS}}\ x_{n}\rrbracket_{G}(T)=$ $\bigcup_{\beta\in T}\{(x_{1}\mapsto\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G},\ldots,x_{n}\mapsto\llbracket\mathcal{E}_{n}\rrbracket^{\beta}_{G})\}$ (57)
@@ 6 Next steps towards the MPG vision
In this paper we have highlighted the critical need to break down the barriers between data and metadata in property graph management. As first concrete steps towards this vision, we introduced a fully specified data model and query language for meta-property graphs, enabling the seamless modeling and interoperation of data and metadata. Our proposals are fully backwards compatible with the ISO standards for property graphs.
While this work lays the foundation for flexible management of property graphs for the requirements of contemporary applications, further research is needed to fully realize this vision. Next steps include investigating: (1) physical representations and indexing strategies for MPGs; (2) query evaluation and optimization solutions for the novel capabilities of MetaGPML; (3) extending the GQL and SQL/PGQ languages with MetaGPML; (4) schema and constraint languages for MPG, building upon the PG-SCHEMA framework [2]; and, (5) educational and training resources for students and professionals in working with MPGs and MetaGPML.

### v2

@@ 5.2 Semantics
Conditions: $\llbracket\mathcal{E}_{1}=\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ $=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (49) $\llbracket\mathcal{E}_{1}<\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ $=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}<\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (50) $\llbracket\text{{x:$\ell$}}\rrbracket^{\beta}_{G}$ $=\begin{cases}True&\text{if }\llbracket\text{{x}}\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\in\mu(\lambda(\llbracket\text{{x}}\rrbracket^{\beta}_{G}))\\ False&\text{if }\llbracket\text{{x}}\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\notin\mu(\lambda(\llbracket\text{{x}}\rrbracket^{\beta}_{G}))\end{cases}$ (51)
Logical operations: $\llbracket\Phi_{1}\ \text{{AND}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\wedge\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (52) $\llbracket\Phi_{1}\ \text{{OR}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\vee\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (53) $\llbracket\text{{NOT}}\ \Phi\rrbracket^{\beta}_{G}$ $=\neg\llbracket\Phi\rrbracket^{\beta}_{G}$ (54)
Set operation: $\llbracket c\ \text{{ELEMENTOF}}\ \text{{y}}\rrbracket^{\beta}_{G}$ $=\begin{cases}True&c\in\mu(\llbracket\text{{y}}\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (55) $\llbracket\text{{SUBSETEQ(x, y)}}\rrbracket^{\beta}_{G}$ $=\begin{cases}True&\mu(\llbracket\text{{x}}\rrbracket^{\beta}_{G})\subseteq\mu(\llbracket\text{{y}}\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (56)
@@
Clauses and queries: $\llbracket\text{{MATCH}}\ \Pi\rrbracket_{G}(T)=\bigcup_{\beta\in T}\{\beta\Join\beta^{\prime}\mid\beta^{\prime}\in\llbracket\Pi\rrbracket_{G},\beta\sim\beta^{\prime}\}$ (57) $\llbracket\text{{FILTER}}\ \Phi\rrbracket_{G}(T)=\{\beta\in T\mid\llbracket\Phi\rrbracket^{\beta}_{G}=True\}$ (58) $\llbracket\mathcal{C}\ \mathcal{Q}\rrbracket_{G}(T)=\llbracket\mathcal{Q}\rrbracket_{G}(\llbracket\mathcal{C}\rrbracket_{G}(T))$ (59) $\llbracket\text{{RETURN}}\ \mathcal{E}_{1}\ \text{{AS}}\ x_{1},\ldots,\mathcal{E}_{n}\ \text{{AS}}\ x_{n}\rrbracket_{G}(T)=\bigcup_{\beta\in T}\{(x_{1}\mapsto\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G},\ldots,x_{n}\mapsto\llbracket\mathcal{E}_{n}\rrbracket^{\beta}_{G})\}$ (60)
@@ 6 Next steps towards the MPG vision
In this paper, we highlighted the critical need to break down barriers between data and metadata in property graph management. As the first concrete steps towards this vision, we introduced a fully specified data model and query language for meta-property graphs, enabling seamless modeling and interoperation of data and metadata. While this work lays the foundation for flexible property graph management for contemporary applications, further research is needed to fully realize this vision.
(1) Physical implementation and technical challenges.
A key challenge is implementing the MPG data model efficiently. Since label sets and properties are treated as data objects with identifiers, they may need dedicated storage and indexing strategies. Alternatively, approaches like concatenated IDs could maintain existing storage structures but may impact query evaluation. Research is needed on physical representations and indexing strategies that optimize MPG performance.
(2) Meta-Property Graphs in practice.
We need to investigate how Meta-Property Graph can enhance knowledge engineering and management in practice. Understanding how metadata awareness and substructure reification can contribute to improving tasks such as auditing and human-in-the-loop validation of knowledge graphs or data cleaning, wrangling, integration, and exchange is crucial. Furthermore, MPGs can facilitate knowledge reasoning and analysis on knowledge graphs by introducing inherent subgraph annotation and querying different forms of metadata. Additionally, developing effective educational approaches and training resources for students and professionals working with MPGs and MetaGPML requires further study.
(3) Improvements and integration.
Meta-Property Graph and MetaGPML can be enhanced through: (1) extending MetaGPML with additional functions to better leverage metadata awareness and existing property graph capabilities like paths, (2) developing schema and constraint languages for MPG building on PG-SCHEMA [pgschema], and (3) incorporating other forms of metadata like reflection to expand the metadata awareness in MPG.
Appendix A Appendix
This appendix provides an in-depth exploration of the semantics missing from section 5.2, supplementing the concise overview presented in the main body of the paper. For those interested in a more detailed understanding, the following sections offer comprehensive insights into the formal structures and rules that underpin the query language’s design and functionality. This additional material aims to enhance the discussion’s clarity and completeness without disrupting the main text’s narrative flow.
$\llbracket\text{{(:$\ell$)}}\rrbracket_{G}=\{()\mid n\in N,\ell\in\mu(\lambda(n))\}$ (61)
$\llbracket\text{{(::$\pi$)}}\rrbracket_{G}=\llbracket\text{{()}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (62)
$\llbracket\text{{(:$\ell$::$\pi$)}}\rrbracket_{G}=\llbracket\text{{(:$\ell$)}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (63)
$\llbracket\text{{(x::$\pi$)}}\rrbracket_{G}=\llbracket\text{{(x)}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (64)
$\llbracket\text{{(:?y)}}\rrbracket_{G}=\{(\text{{y}}\mapsto l)\mid n\in N,l=\lambda(n)\}$ (65)
$\llbracket\text{{(:?y::$\pi$)}}\rrbracket_{G}=\llbracket\text{{(:?y)}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (66)
$\llbracket\text{{(x:?y::$\pi$)}}\rrbracket_{G}=\llbracket\text{{(x:?y)}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (67)
$\llbracket\text{{().z}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid n\in N,p\in\sigma(n)\}$ (68)
$\llbracket\text{{(:$\ell$).z}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid n\in N,\ell\in\mu(\lambda(n)),p\in\sigma(n)\}$ (69)
$\llbracket\text{{(::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{().z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (70)
$\llbracket\text{{(x:$\ell$).z}}\rrbracket_{G}=\{(\text{{x}}\mapsto n,\text{{z}}\mapsto p)\mid n\in N,\ell\in\mu(\lambda(n)),p\in\sigma(n)\}$ (71)
$\llbracket\text{{(:$\ell$::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{(:$\ell$).z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (72)
$\llbracket\text{{(x::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{(x).z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (73)
$\llbracket\text{{(x:$\ell$::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{(x:$\ell$).z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (74)
$\llbracket\text{{(:?y).z}}\rrbracket_{G}=\{(\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid n\in N,l=\lambda(n),p\in\sigma(n)\}$ (75)
$\llbracket\text{{(x:?y).z}}\rrbracket_{G}=\{(\text{{x}}\mapsto n,\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid n\in N,l=\lambda(n),p\in\sigma(n)\}$ (76)
$\llbracket\text{{(:?y::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{(:?y).z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (77)
$\llbracket\text{{(x:?y::$\pi$).z}}\rrbracket_{G}=\llbracket\text{{(x:?y).z}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (78)
-$\llbracket\text{{-[]-}}\rrbracket_{G}=\{()\mid e\in E_{u}\}$ (79)
-$\llbracket\text{{-[x]-}}\rrbracket_{G}=\{(\text{{x}}\mapsto e)\mid e\in E_{u}\}$ (80)
-$\llbracket\text{{-[:$\ell$]-}}\rrbracket_{G}=\{()\mid e\in E_{u},\ell\in\mu(\lambda(e))\}$ (81)
-$\llbracket\text{{-[:?y]-}}\rrbracket_{G}=\{(\text{{y}}\mapsto l)\mid e\in E_{u},l=\lambda(e)\}$ (82)
-$\llbracket\text{{-[x:?y]-}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{y}}\mapsto l)\mid e\in E_{u},l=\lambda(e)\}$ (83)
-$\llbracket\text{{-[].z-}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid e\in E_{u},p\in\sigma(e)\}$ (84)
-$\llbracket\text{{-[x].z-}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{z}}\mapsto p)\mid e\in E_{u},p\in\sigma(e)\}$ (85)
-$\llbracket\text{{-[:$\ell$].z-}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid e\in E_{u},\ell\in\mu(\lambda(e)),p\in\sigma(e)\}$ (86)
-$\llbracket\text{{-[x:$\ell$].z-}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{z}}\mapsto p)\mid e\in E_{u},\ell\in\mu(\lambda(e)),p\in\sigma(e)\}$ (87)
-$\llbracket\text{{-[:?y].z-}}\rrbracket_{G}=\{(\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid e\in E_{u},l=\lambda(e),p\in\sigma(e)\}$ (88)
-$\llbracket\text{{-[x:?y].z-}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid e\in E_{u},l=\lambda(e),p\in\sigma(e)\}$ (89)
-$\llbracket\text{{-[x:$\ell$]->}}\rrbracket_{G}=\{(\text{{x}}\mapsto e)\mid e\in E_{d},\ell\in\mu(\lambda(e))\}$ (90)
-$\llbracket\text{{-[:?y]->}}\rrbracket_{G}=\{(\text{{y}}\mapsto l)\mid e\in E_{d},l=\lambda(e)\}$ (91)
-$\llbracket\text{{-[].z->}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid e\in E_{d},p\in\sigma(e)\}$ (92)
-$\llbracket\text{{-[:$\ell$].z->}}\rrbracket_{G}=\{(\text{{z}}\mapsto p)\mid e\in E_{d},\ell\in\mu(\lambda(e)),p\in\sigma(e)\}$ (93)
-$\llbracket\text{{-[x:$\ell$].z->}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{z}}\mapsto p)\mid e\in E_{d},\ell\in\mu(\lambda(e)),p\in\sigma(e)\}$ (94)
-$\llbracket\text{{-[:?y].z->}}\rrbracket_{G}=\{(\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid e\in E_{d},l=\lambda(e),p\in\sigma(e)\}$ (95)
-$\llbracket\text{{-[x:?y].z->}}\rrbracket_{G}=\{(\text{{x}}\mapsto e,\text{{y}}\mapsto l,\text{{z}}\mapsto p)\mid e\in E_{d},l=\lambda(e),p\in\sigma(e)\}$ (96)